OpenStudy (anonymous):

choose the ordered pair (x,y) that is a solution for the equation. 13x-11y=14 (-6,-3) (-4,-6) (0,-3) (-16,-5)

OpenStudy (anonymous):

Did you try plugging each solution in for x,y and seeing if the sides were equal?

OpenStudy (anonymous):

i wasn't taught how to do that i'm now learning this

OpenStudy (anonymous):

oh ok, Well let's try to put -6 in for x and then -3 for y..... so it would look like this... 13(-6)-11(-3)=14 so 13 time -6 is -78 and 11 times -3 is -33 -78 minus a -33 (so -78 +33) = 45.... 45 is not = to 14... so it's one of the other 3... does that make sense or not? It is so hard to explain online haha :)

OpenStudy (anonymous):

hold on, let me try. that was a good explanation

OpenStudy (anonymous):

If you know PEMDAS (the order of operations) once you know how to plug in the possible answers you can figure out which answer is best. Now solving for x, y on your own is another story. Finding x or y requires you to isolate the x or y by moving parts of the equation to the other side, leaving the x, y on one side. For example If 3x + 10 = 19 then.... 3x = 19 -10 3x = 9 x = 9/3 x = 3

OpenStudy (anonymous):

(-4,-6) right?

OpenStudy (anonymous):

Yep! :)